cannot seem to get this!!! from my teacher's own problem solving packet. it reads...
The larger of two consecutive even integers is six less than twice the smaller. Find the numbers.
thank you tons!!!

Found 3 solutions by mukhopadhyay, venugopalramana, sarah_adam:
Answer by mukhopadhyay(490)

If one of the even integers is x, the other is x+2;
Per question, (x+2) = 2x-6
=> x = 8;
The numbers are 8 and 10.

Answer by venugopalramana(3286)

The larger of two consecutive even integers is six less than twice the smaller. Find the numbers.
thank you tons!!!
LET SMALLER EVEN NUMBER BE =2N...
TWICE THIS =2*2N=4N....6 LESS =4N-6...THIS IS EQUAL TO NEXT EVEN NUMBER =2N+2
4N-6=2N+2
4N-2N=2+6=8
2N=8
N=4....SO THE 2 NUMBERS ARE 2*4=8 AND 10.

Answer by sarah_adam(201)

The two consequtive even integers are n ,(n+2)
greater integer is (n+2) is six less then twice the smaller one
so (n+2) = 2n - 6
2n - n = 6 +2
n = 8
n+2 = 10
the two integers are 8,10